Halliford Corporation expects to have earnings this coming year of $ 2.85 per share. Halliford plans to retain all of its earnings for the next two years. For the subsequent two​ years, the firm will retain 47 % of its earnings. It will then retain 23 % of its earnings from that point onward. Each​ year, retained earnings will be invested in new projects with an expected return of 20.33 % per year. Any earnings that are not retained will be paid out as dividends. Assume​ Halliford's share count remains constant and all earnings growth comes from the investment of retained earnings. If​ Halliford's equity cost of capital is 8.5 %​, what price would you estimate for Halliford​ stock? ​Note: Remenber that growth rate is computed​ as: retention rate times rate of return.
Answer #1
1) The first step is to find the growth rate:
Growth rate = ROE*b
where ROE = return on equity
b = retention ratio
Growth rate for the 2nd and 3rd years = 20.33%*100% = 20.33%
Growth rate for years 4 and 5 = 20.33%*0.47) = 9.56%
Perpetual constant growth rate from 65th year = 20.33%*0.23 4.68%
2) Expected EPS, dividends and their PV:
Year EPS [$] Growth rate [%] Payout [%] Dividend [$] PVIF at 8.5% PV at 8.58%
1 2.85 0 0 0.92166 $                 -  
2 3.429 20.33 0 0 0.84946 $                 -  
3 4.127 20.33 53% 2.19 0.78291 $            1.71
4 4.521 9.56 53% 2.40 0.72157 $            1.73
5 4.953 9.56 77% 3.81 0.66504 $            2.53
6 onwards 5.185 4.68 77% 3.99
3) Estimated price of the stock PV of dividends for years 3 to 5 $ 5.97
Terminal value of dividends 3.99/(0.085-0.0468) = $ 104.45
PV of terminal value = 104.45*0.665 = $ 69.46
Estimated price of the stock= 5.97+69.46 = $ 75.43